1 Conduct of Operations for Hall C 2 E12-06-110, E12-06-121 3 February 7, 2018

4 Contents

1 Preface  2

5

2 Documentation  2

6

3 Shift Personnel Training  3

7

4 Organization and Administration  4

8

4.1 Run Coordinator . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

9

4.2 Physics Division Liaison . . . . . . . . . . . . . . . . . . . . . . . 6

10

4.3 Hall Work Coordinator . . . . . . . . . . . . . . . . . . . . . . . . 6

11

4.4 Shift Leader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

12

4.5 Shift Member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

13

4.6 Accelerator Operations Hall Liaison . . . . . . . . . . . . . . . . 8

14

4.7 Accelerator Physicist Liaison . . . . . . . . . . . . . . . . . . . . 9

15

4.8 Engineering Liaison . . . . . . . . . . . . . . . . . . . . . . . . . 9

16

5 Operating Procedures  9

17

5.1 Shift Routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

18

5.1.1 Operating Shifts . . . . . . . . . . . . . . . . . . . . . . . 9

19

5.1.2 Standby Shifts . . . . . . . . . . . . . . . . . . . . . . . . 10

20

5.1.3 Operations Turnover . . . . . . . . . . . . . . . . . . . . . 10

21

5.1.4 Timely Orders to Operators . . . . . . . . . . . . . . . . . 10

22

5.1.5 Operator Aid Postings . . . . . . . . . . . . . . . . . . . . 10

23

5.2 Hall Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

24

5.3 Collaboration Request for Laboratory Resources . . . . . . . . . 12

25

5.4 Scheduling of Work by Outside Groups . . . . . . . . . . . . . . 12

26

5.5 Control of Equipment and System Status . . . . . . . . . . . . . 12

27

5.6 Equipment Labeling . . . . . . . . . . . . . . . . . . . . . . . . . 13

28

5.7 Independent Verification . . . . . . . . . . . . . . . . . . . . . . . 13

29

5.8 Logkeeping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

30

A Special Procedures for Hall C  14

31

B Special Procedures for E12-06-110, E12-06-121  15

32

C Signature Sheets  16

33

34 1 Preface

35 As part of its mission, JLab provides the resources necessary for international 36 collaborations of scientists to carry out basic research in nuclear physics and 37 related disciplines. This research must be conducted in a manner that ensures 38 that environmental, health and safety (EH&S) concerns receive the highest con-39 sideration. At the same time the programmatic goals of the laboratory require 40 that it produce the highest quality physics results efficiently. 41 Guidance on how to balance thoughtful, measured EH&S concerns with 42 efficient operation has been taken from the Director’s Safety Council, the JLab 43 EH&S Manual, and the JLab Director’s Office. A graded approach is followed 44 in which the measures taken are matched to the scale, cost, complexity, and 45 hazards of the operation. 46 This document outlines how approved experiment collaborations 47 will conduct operations in a safe and effective manner during the time 48 period that E12-06-110, E12-06-121 are on the floor. Installation pe-49 riods are not covered by this document. Furthermore, this document 50 is directed to physics users and physics staff rather than the Hall 51 C technical staff. It must be read, understood, and followed by all 52 members of the collaboration.

53 2 Documentation

54 This experiment uses the standard Hall C equipment, the Moller polarimeter, 55 and a polarized ^{3}He target. All of the procedures to be used during the course 56 of the experiment are contained in the following documents^{1}:

57 • The Conduct of Operations for Hall C E12-06-110, E12-06-121 (COO), 58 the document you are now reading.

59 • Experiment Safety Assessment Document (ESAD) for E12-06-110, E12-60 06-121 (referring to the base equipment as well as any experiment-specific 61 changes)

62 • Radiation Safety Assessment Document (RSAD)

63 • JLab Emergency Response Guidelines (ERG)

64 • Hall C Standard Equipment Manual

65 Reference copies of these documents will be available in the Counting House 66 for the duration of the experiment. The present document shall hereafter be 67 referred to as the COO. The Experiment Safety Assessment Document shall 68 hereafter be referred to as the ESAD, and the Radiation Safety Assessment 69 Document shall be referred to as the RSAD. The ESAD and COO may also be 70 available on the WWW at an experiment-specific web site^{2}. The COO, the 71 ESAD and the RSAD are required reading for shift personnel. 72 A full description of the physics motivation for the experiments, collabora-73 tion lists, and general plans for carrying out an experiment can be found in the 74 proposal(s) to the JLab Program Advisory Committee (PAC).

1The process is documented at http://www.jlab.org/user_resources/PFX

75 3 Shift Personnel Training

76 All personnel on shift are required to have successfully completed and be current 77 in the following JLab safety training:

78 • EH&S Orientation (SAF 100)

79 • Radiation Worker Training (SAF 801)

80 • Oxygen Deficiency Hazard Training (SAF 103)

81 • Hall C Safety Awareness Walk-Through ( SAF112 )

82 • Conduct of Operations (SAF 120)

83 All experiment personnel are required to have radiation badges in their pos-84 session during their shifts. The Safety Awareness Walk-Through will emphasize 85 hazards that are typical of normal Hall operations. Hazards peculiar to the 86 current experimental setup are addressed in the appendices of this document. 87 In addition, all shift personnel will be trained in the safety procedures to be 88 followed for access to the Hall during restricted or contolled access periods (see 89 Appendix B for special requirements during the “laser lockdown” period). The 90 Hall access training will include a brief discussion of the purpose and opera-91 tion of the Personnel Safety System (PSS) for the Hall. Individuals within the 92 collaboration may be required to have other equipment or procedure-specific 93 training. The need for such training shall be determined by the experiment 94 spokesperson in consultation with the Hall Leader and Physics Division Safety 95 Officer. 96 In addition, experiment personnel must familiarize themselves with the sec-97 tions of the JLab EH&S Manual relevant for their work in the Hall. A reference 98 copy of this document is available in the main hallway of the Counting House.

99 It is also available via http://www.jlab.org/ehs/ehsmanual/index.html.

100 Finally, JLab Lock and Tag^{3} training is required for all staff/users who will 101 be performing maintenance on electrical and mechanical equipment which can-102 not be physically and verifiably isolated from an energy source. This training, 103 SAF104, can be found at:

104 http://www.jlab.org/div_dept/train/webbasedtraining.html.

2https://hallcweb.jlab.org/experiments/A1n 3The EH&S Manual provides Lockout/Tagout information in Chapter 6110 at http://www.jlab.org/ehs/ehsmanual/manual6110.html.

Figure 1: Functional Organization of the Hall C Team. Dashed lines indicate information flow, solid lines indicate responsibility.

105 4 Organization and Administration

106 The operation of the experiment is directed by the Spokespersons and the Hall 107 Leader, Thia Keppel. An organization chart for the experiment is found in 108 Figure 1.

109 4.1 Run Coordinator

110 The Run Coordinator is the immediate on-site manager of the experiment and 111 is responsible for ensuring that the physics goals of the experiment are met. 112 This individual is designated by the experiment spokespersons and approved 113 by the Hall Leader. The Run Coordinator shall ensure that the Hall Group 114 Leader, Physics Division Liaison, and at least one Spokesperson are aware of all 115 pertinent issues. The Run Coordinator shall promote an environment in which 116 the highest safety standards are maintained. All Run Coordinators must 117 ensure that all of the JLab training necessary to perform their duties 118 is up to date before their shift as Run Coordinator commences. The 119 functions of the Run Coordinator are: 120 I. To manage daily operation of the experiment:

121 • to ensure that the run plan is clear to the shift workers.

122 • to define the data quality appropriate for the goals of each shift.

123 • to track the progress of the experiment.

124 • to coordinate and schedule activities (e.g., Hall accesses) in order to opti-125 mize productivity.

126 • to ensure that an experiment checklist is completed every 24 hrs during 127 standby shifts.

128 • together with the Physics Division Liaison, to ensure that the counting 129 house is manned appropriately: i.e., sufficient personnel are present to 130 safely carry out the experimental program or monitor the apparatus as 131 needed.

132 II. To coordinate interactions between JLab and the experiment. This entails:

133 • to ensure that the Hall C Group Leader and Experiment Spokespeople are 134 aware of all necessary issues.

135 • informing the Program Deputy of the experiment’s status and plans at a 136 7:45 AM program deputy/halls meeting in the MCC during the working 137 week, and at an agreed upon time on weekends or holidays.

138 • representing the collaboration at the 8:00 AM daily summary meeting in 139 the MCC during the work week.

140 • attending the 1:30 PM Wednesday scheduling meeting in the MCC con-141 ference room to represent the collaboration and to present a report on the 142 preceding week.

143 • remaining in the local area and being available by cell-phone/pager at all 144 times. (If temporarily unavailable the Run Coordinator must designate 145 another qualified collaborator as a replacement.)

146 • in conjunction with the Hall Work Coordinator, scheduling work by groups 147 outside the collaboration.

148 • interact with the Accelerator Program Deputy to plan and conduct un-149 scheduled activities.

150 • in conjunction with the Hall Work Coordinator, scheduling work by groups 151 outside the collaboration. This work will normally coincide with the 152 scheduled machine maintenance days. This coordination requires a weekly 153 meeting of these two individuals. The product of this meeting will include 154 any necessary updates to the “Access Authorization List”.

155 • to be responsible for safe transition of the Hall to Restricted Access in 156 coordination with the Hall work coordinator.

157 • to provide an oral report at the weekly Hall C meeting^{4} updating the 158 experimental progress to the collaboration.

159 III. To submit a written report to the Hall Leader which includes run time 160 statistics and a description of any significant problems with the Hall instrumen-161 tation.

162 4.2 Physics Division Liaison

163 Broadly speaking, the Physics Division Liaison to the experiment is a Hall C 164 staff member selected by Thia Keppel to oversee the hall’s interests with re-165 spect to personnel and equipment protection.^{5} This is true for all four halls. 166 However, the role of the Physics Division Liaison may include other responsibil-167 ities depending upon the experiment and other factors. His/her responsibilities 168 include:

169 • Oversee that proper rules of safety are carefully followed in the conduct 170 of the experiment.

171 • Approve a Hall status change to Restricted Access in coordination with 172 the Hall Work Coordinator.

173 • Training verification of shift workers via JList software.

174 • Together with the Run Coordinator, ensure that the counting house is 175 manned appropriately: i.e., sufficient personnel are present to safely carry 176 out the experimental program or monitor the apparatus as needed.

177 4.3 Hall Work Coordinator

178 The Hall Work Coordinator’s responsibilities are:

179 • to act as the single point of contact for all work in the hall.

4typically held at 3:00pm on Monday.

5The responsibilities described here correspond to those of the Physics Division Liaison during the operating phase of the experiment as outlined in the EH&S Manual Chapter 3120/Glossary.

180 • to determine if the scheduled activities in the hall can be done safely. 181 These activities shall be coordinated with the Physics Division Liaison 182 and the Run Coordinator. Tasks should also be inputted into the work

183  task lists http://www.jlab.org/listsites/.

184 • to ensure that workers are properly trained, are familiar with all significant 185 hazards, and are aware of all applicable work control documents associated 186 with the project.

187 • in coordination with the Physics Division Liaison, ensure that the hall 188 apparatus is made safe before giving permission to make a transition to 189 Restricted Access (e.g., turn off unused magnets, install protective shields 190 as needed, fulfill specific requirements in the ESAD, etc.).

191 4.4 Shift Leader

192 Each shift is led by a Shift Leader. The selection of shift leaders is the respon-193 sibility of the Run Coordinator and Physics Division Liaison. The Shift Leader 194 has the following responsibilities:

195 • to carry out the scientific program planned for the shift in a safe and 196 efficient manner.

197 • to ensure that the logbook contains a complete and accurate description 198 of the events and actions which occurred during the shift.

199 • to serve as primary contact between the machine control center (MCC) 200 and experiment personnel.

201 • to oversee that hall equipment is operated properly.

202 • to ensure the shift checklist is performed every eight hours on operating 203 shifts.

204 • to ensure that equipment malfunctions are properly labeled and locked-205 out if necessary and to communicate this to shift personnel and subsystem 206 experts.

207 • to note in the logbook when workers from outside groups (such as survey 208 and alignment) stop by the counting house before entering the hall when 209 in Controlled Access. Furthermore, to confirm that these workers have 210 communicated with the Run Coordinator and the Hall Work Coordinator.

211 • to coordinate the response of the shift crew to any emergency situation, 212 including the notification of appropriate individuals as outlined in the Hall 213 C Emergency Response Guidelines (ERG).

214 • to ensure that in any emergency situation the experiment Physics Division 215 Liaison, Run Coordinator, and Hall Leader are notified immediately.

216 • to notify the Run Coordinator and the Hall Leader, if the hall is down 217 due to equipment failure for more than four hours.

218 The Shift Leader has the following authority:

219 • to assign tasks to the shift members as needed.

220 • to request that the state of the hall be changed (Request for a change to 221 Restricted Access must be approved by the Physics Division Liaison.)

222 • to limit the number of people in the Counting House or hall if required to 223 effectively and safely carry out the experiment.

224 • to limit access to hall on-line computers if required to effectively and safely 225 carry out the experiment.

226 • to authorize qualified personnel to make modifications in the experiment 227 configuration within the allowed parameters, as specified in the standard 228 equipment manual.

229 • to authorize time accounting for the shift.

230 4.5 Shift Member

231 The responsibilities of each shift member are to:

232 • carry out the scientific goals of the shift in a safe and efficient manner 233 under direction of the shift leader.

234 • read the logbook to be aware of changes in goals, operating parameters, 235 and new documentation.

236 • monitor the equipment for problems.

237 • maintain adequate records of the progress of the shift.

238 • be present before the start of each shift and coordinate current operating 239 conditions with the previous shift.

240 • keep all training up-to-date.

241 4.6 Accelerator Operations Hall Liaison

242 Each physics hall has an Accelerator Operator or Crew Chief assigned as a Hall 243 Liaison. The Hall Liaison helps to facilitate information exchange between the 244 experimenters and the MCC Operations Group, both in advance of and during 245 actual experiments. The Hall Liaison, among other things, is responsible for 246 making sure that experiment-specific information, procedures and requirements 247 are available to all other operators and Crew Chiefs so that beam delivery can 248 proceed efficiently. The Hall C liaison is Daniel Moser.

249 4.7 Accelerator Physicist Liaison

250 The Accelerator Physicists Experiment Liaison serves as the primary contact 251 on hall beam physics issues for the Physics, Accelerator and Engineering Divi-252 sions. This liaison owns the process of establishing physics quality beam to the 253 experiment including developing beam optics configurations capable of meet-254 ing the experiments requirements, identifying tools needed to diagnose, monitor 255 and verify beam performance during the experiment as well as developing beam 256 startup, setup and commissioning plans. The Hall C liaison is Jay Benesch.

257 4.8 Engineering Liaison

258 Each experiment conducted at JLab will be evaluated to determine if its com-259 plexity requires facilitation with the Engineering Division to help ensure a suc-260 cessful outcome. Experiments that require facilitation will be assigned an in-261 dividual from the Engineering Division to act as liaison between the Division 262 and the associated Experimental and Physics Division staff. The liaison acts as 263 a single point contact in order to facilitate information exchange between the 264 experimenters and those in the Engineering Division responsible for, but not 265 limited to, the systems requirements, design, scheduling, fabrication, installa-266 tion, testing, documentation, and budgeting. Ideally, the liaison is aware of all 267 work conducted by Engineering for the experiment and ensures the appropriate 268 resources are defined and allocated. Any issues and/or concerns are identified, 269 documented, and tracked. 270 For the current run period, the review found that no such liaison was re-271 quired.

272 5 Operating Procedures

273 5.1 Shift Routines

274 There are two types of shifts for active hall experiments: Operating and Standby. 275 Operating shifts are the normal status when beam is available for the exper-276 iment. Standby shifts are periods designated by the Run Coordinator when 277 beam is not available or not in use in the hall and none of the equipment, ex-278 cept for the target, requires continuous monitoring. Standby status may result 279 from normal operational planning or from abnormal conditions such as a major 280 down time due to equipment failure.

281 5.1.1 Operating Shifts

282 During operating shifts, 24 hour occupation of the counting house area will be 283 maintained by crews of at least two persons ^{6} in 8 hour shifts. One person per 284 shift is designated as the Shift Leader.

6The readiness review committee may require more personnel depending on the complexity of the experiment. Two people are the minimum required for safe operations.

285 The number of persons assigned to a shift will depend on the tasks assigned 286 during the shift. A shift schedule will be posted in the Counting House listing 287 the times and names of personnel on shift and identifying the Shift Leader and 288 Run Coordinator, cell 757-270-8916. The shift schedule may be available at 289 an experiment-specific website. The Run Coordinator may also designate and 290 supervise other teams for duties such as offline analysis.

291 5.1.2 Standby Shifts

292 During Standby shifts, shift personnel are not required to be on site at JLab 293 but must be available through telephone contact to come in if they are needed. 294 Monitoring the target system can require the presence of a Target Operator 295 during a standby shift. The Target Operator then also acts as Shift Leader. 296 The Run Coordinator will ensure that the shift checklist is executed at least 297 once every 24 hours.

298 5.1.3 Operations Turnover

299 The electronic log book, accessible from the web, is a very effective means of 300 remotely obtaining information about experimental operations. This allows ex-301 perimenters to log in remotely and view all log book entries prior to commencing 302 their shift. Information which can only be recorded in the paper log book, should 303 be noted accordingly, point to in the electronic logbook, and communicated be-304 tween incoming and outgoing shift personnel directly. 305 Efficient and effective shift changeovers during experiment operation are 306 enhanced by overlapping shifts. Therefore, whenever possible, shift leaders and 307 workers are scheduled in shifts that are staggered by four hours, leading to an 308 overlap of half a shift. If this is not the case, shift members must show up 309 ten minutes prior to shift start (and plan to stay ten minutes after) for the 310 purpose of information exchange to those taking over the same tasks. In all 311 cases incoming shift leaders must discuss the experiment and Hall status with 312 the outgoing shift leaders.

313 5.1.4 Timely Orders to Operators

314 The initial run plan is the responsibility of the Run Coordinator and shall be 315 clearly recorded in the log book. This plan specifies the tasks to be performed 316 in the next 48–72 hours, including any special conditions or data runs, updated 317 documentation and its location and/or alternate plans. Any changes to the run 318 plan shall be recorded in the log book and the white board in the counting 319 house.

320 5.1.5 Operator Aid Postings

321 The day-to-day schedule, contact instructions for key personnel, and any other 322 information relevant to current activities are located on the white board in the 323 Counting House. Shift personnel should consult the white board, especially 324 at the beginning of their shift, to be aware of any updates to current running 325 conditions. 326 Information pertaining to daily activities in Hall C must be posted on the 327 bulletin board or written on the white board at the entrance to the hall.

328 5.2 Hall Access

329 Work in designated radiation areas will be carried out in accordance with the 330 JLab RadCon Manual. In particular, no material may be removed from the hall 331 after beam delivery without proper approval from the RadCon Group. During 332 operations, no one is allowed in the hall without either being accompanied, or 333 informing shift personnel and checking in on a regular basis. 334 During a running experiment the hall will normally be in Beam Permit. 335 When temporary access to the hall is needed the Shift Leader can ask the MCC 336 to bring the hall to Controlled Access. If long term access to the hall is required, 337 the Shift Leader may request the hall be brought to Restricted Access. Such 338 a request requires prior approval from the Physics Division Liaison, while the 339 actual transition will be supervised by the Hall Work Coordinator. 340 Restricted Access is a state where delivery of beam and/or RF power is not 341 permitted, and entry to and exit from the hall is not controlled by the Personnel 342 Safety System. This is the normal state of the hall when the accelerator is off 343 and no experiments are running. Access is “restricted” only in the sense that 344 the hall is not open to the general public. Well-defined check-list procedures are 345 to be followed whenever the hall is brought to and from Restricted Access. 346 Restricted Access is the period when all major work must be completed in 347 the hall. Consequently, all activities require advanced planning and must be 348 scheduled for resources and safe operation. In order to streamline the activities 349 in the hall and ensure everyone has ready access to the current status and 350 requirements for work, there are two important resources:

351 • Single point of contact, which is the “Hall Work Coordinator”

352 • Information board at the entrance to the hall

353 All work must be scheduled through the Hall Work Coordinator. The content 354 on the information board is the responsibility of the hall safety wardens and 355 the Hall Work Coordinator. The information board will contain all critical in-356 formation required for safe entry into the hall. This information will include 357 a succinct, one page safety summary covering the hall’s current safety hazards 358 and mitigating measures (to be read by all persons working in the hall), ac-359 tive Operational Safety Procedures (OSPs) and Temporary Operational Safety 360 Procedures (TOSPs), required temporary work permits (e.g., Radiation Work 361 Permits), current activities in the hall, points of contact, and required training 362 and safety equipment.

363 5.3 Collaboration Request for Laboratory Resources

364 The collaboration may request additional services from Accelerator Division 365 through the Accelerator Division Liaison, Geoffrey Krafft. Alternatively, the 366 collaboration may also request additional services from hall personnel through 367 the Physics Division Liaison, Greg Smith. These requests should be noted in 368 the logbook. Some requests may require that an OSP, or TOSP be developed. 369 Major, abnormal, or unanticipated configuration modifications such as stack-370 ing or movement of significant shielding, unanticipated vacuum work, unantici-371 pated beam line modifications, the replacement of a wire chamber, etc., require 372 approval of the Hall C Leader, Thia Keppel ^{7}, and the use of appropriate per-373 sonnel. The Hall Leader may require that a OSP, or TOSP be prepared.

374 5.4 Scheduling of Work by Outside Groups

375 Work in the hall that is to be performed by groups outside the collaboration 376 such as survey and alignment, plant services, air conditioning, etc., must be 377 scheduled so that it does not endanger personnel or equipment or interfere with 378 the experiment. Non-emergency activities by these groups should be scheduled 379 to coincide with the planned accelerator maintenance periods. To maximize 380 efficiency, the Run Coordinator (representing the collaboration) and the Hall 381 Work Coordinator (representing Hall C) will concur on work scheduling. The 382 Hall Work Coordinator’s job is to coordinate activities in the hall so that work 383 can take place smoothly and safely and to insure that multiple activities do not 384 interfere. 385 The Work Coordinator and the Run Coordinator will meet as needed to 386 plan the work scheduled for the upcoming maintenance period. The product 387 of this meeting will be a list of work in the hall, the required access state of 388 the hall (Controlled or Restricted), appropriate work control documents, and 389 educational or other safety measures (such as escorts) that are needed. 390 The ATLis should be used for coordinating the cross divisional work activi-

391 ties http://www.jlab.org/listsites/.

392 5.5 Control of Equipment and System Status

393 The operation of the standard experimental equipment is documented in the 394 Hall C Standard Equipment Manual. This document includes information on 395 the normal response to alarms and equipment malfunctions. 396 The ESAD and Hall C Standard Equipment Manual lists the authorized 397 subsystem experts. This list may be amended as necessary to reflect personnel 398 and training changes with the authorization of the subsystem expert. A copy 399 of these amendments will be attached to the main document and kept in the 400 Counting House.

7Configuration changes as outlined above can affect site boundary dose and the production of airborne radioactivity. They require consulting with RadCon or EH&S personnel, as appropriate.

401 All general equipment installation, maintenance, and testing activities are 402 to be carried out in accordance with the JLab EH&S Manual.

403 5.6 Equipment Labeling

404 The experiment and hall equipment shall be properly labeled so it can be quickly 405 identified by both shift and maintenance personnel. Proper labeling helps pre-406 vent incorrect operation or modification of equipment by non-experts and fa-407 cilitates proper and efficient operation by qualified personnel. Labeling also 408 increases the likelihood that proper procedures will be followed in case of emer-409 gency. 410 Improper labels should be corrected immediately if possible. Otherwise, the 411 Shift Leader should be notified so that correct labeling can be requested from 412 the qualified expert.

413 5.7 Independent Verification

414 The Run Coordinator will provide the shift crew with a set of measures for 415 checking the quality of the experimental data. The up-to-date Hall C shift 416 checklist (and instructions) shall be made available to shift personnel at hall-417 specific sites on the data acquisition computers. The checklist will be completed 418 at least once per shift during operating shifts and once per day during standby 419 shifts. Additional items may be added to the list by the Run Coordinator or 420 subsystem experts. 421 The Hall C work coordinator provides more general check lists for closing 422 the experimental Hall and conditions when the Hall is used as an accelerator 423 dump.

424 5.8 Logkeeping

425 Shift personnel will update the electronic logbook, which serves as the record of 426 the experiment. The quality of the information recorded in the logbook deter-427 mines the utility of the data. All data recorded electronically will be referenced 428 in the computer logbook with the appropriate run number and run information. 429 All relevant activities are to be recorded, including all changes of experiment 430 conditions and equipment failures. 431 Checklists performed using Hall C-specific forms should also be scanned into 432 the computer logbook when completed. The completed paper forms should be 433 stored in a binder in the counting house. All deviations from normal operating 434 parameters shall be recorded in the logbook. 435 The computer logbook will also serve as the primary reference for the deter-436 mination of the operational efficiency of the experimental apparatus in the Hall. 437 As such it is essential that it provide an accurate record of the capability of the 438 equipment to carry out the intended research program. Finally, the computer 439 logbook is the place of record for all safety issues and introductions of new or 440 updated documentation and procedures.

441 A Special Procedures for Hall C

442 There are no special operating procedures for Hall C.

443 B Special Procedures for E12-06-110, E12-06-

444 121

445 Each shift requires a shift leader and a polarized ^{3}He target operator. A third 446 person on shift is extremely useful, but not required. The shift leader has the 447 standard duties of shift leader to ensure proper data taking, log all activity, and 448 to fill out the Beam Accounting form. The target operator should focus on the 449 operation of the polarized ^{3}He target. Target operator training is arranged by 450 Jian-ping Chen. The shift leader or third shift worker will run the DAQ and 451 online analysis codes. 452 This run period uses the “standard” Hall C equipment, a polarized ^{3}He tar-453 get, and the Moller polarimeter. The use and safety procedures of the standard 454 equipment are documented in the Hall C Standard Equipment Manual available 455 from the Hall C web page. The polarized ^{3}He target will be used in Hall C for 456 the first time. The Moller polarimeter, while used to be part of the Hall C 457 standard equipment during the 6 GeV era, will also be used for the first time 458 after the 12 GeV upgrade. The use and safety procedures of the target and 459 the Moller polarimeter are documented in the TOSP (Target Operation Safety 460 Procedures) and the Moller OSP, respectively, available from the experimental 461 webpage of E12-06-110, E12-06-121. 462 One special requirement of the polarized ^{3}He target is the special “laser 463 lockdown” period. During this period, only personnel with polarized ^{3}He target 464 training will be allowed to access the hall. Details of the “laser lockdown” period 465 and its safety and training requirement are documented in the TOSP.

466 C Signature Sheets

467 After reading this document, as well as TOSP, OSP of the Moller polarimetry, 468 ESAD, RSAD, and ERG, workers need to sign the signature sheet located in 469 the “yellow binder” of the experiment specific documents. This binder can be 470 found in the Hall C counting house and in the MCC.